OpenStudy (anonymous):

how far does a rolling wheel with a 10-inch radius travel in ten revolutions?

OpenStudy (anonymous):

radius = 10 inches circumference = 2xpixradius=2xpix10 = 20pi inches put pi= 3.14 once you get the circumference multiply it by 10 you will get the distance traveled in inches

OpenStudy (anonymous):

@03453660 im getting 628 but my homework says thats wrong

OpenStudy (anonymous):

no check the unit that might be in meters

OpenStudy (anonymous):

20 * 3.14= 62.8 62.8*10=628

OpenStudy (anonymous):

628 inches = 15.95 meters = 52.33 feet = 17.44 yards

OpenStudy (anonymous):

yes 628 inches you r right but check the unit please

OpenStudy (anonymous):

oh ok. its asking for \[\pi inches\]